Mathematics 136 – Advanced Placement Calculus Evening Exam 1 Groundrules September 23, 2009

• The exam will start promptly at 7:30pm and end promptly at 9:00pm. (Exception: some students may have been authorized to receive additional time on exams by the Counseling Center. If you believe you are eligible for this accomodation because of a learning disability, you must arrange to have your situation verified by the Counseling Center and I must be officially notified.)

• The exam will be designed to take about 50 minutes to an hour to complete if you are wellprepared and work steadily. But you will have the extra 30 to 40 minutes to work and check your solutions if you need that much time.

• You will be given a TI-30 scientific calculator for the exam which does NOT have graphing capabilities. (Note: Some of you may have a similar calculator purchased for use in Chemistry courses here. Those are also OK.) So be prepared to answer questions without your personal calculator.

• Important: No use of computers, cell phones, pagers, PDA’s, I-pods, or any other electronic devices will be permitted during the exam. Please leave such devices in your room or put them away in your backpack (making sure cell phones are turned off so they do not ring during the exam and disturb you and your classmates).

• You will be provided with an exam booklet where you will place your solutions. This will have plenty of room for scratch work, so it should not be necessary to bring any paper. If you do bring other paper for scratch work, it must be completely blank. I will inspect any paper you bring to the exam to make sure this rule is followed.